# Aladin Djuhera1, Swanand Kadhe2, Farhan Ahmed2, Syed Zawad2, Fernando Koch3, Walid Saad4, Holger Boche1

### v2

@@ Abstract
Fine-tuning large language models (LLMs) on telecom datasets is a common practice to adapt general-purpose models to the telecom domain. However, little attention has been paid to how this process may compromise model safety. Recent research has shown that even benign fine-tuning can degrade the safety alignment of LLMs, causing them to respond to harmful or unethical user queries. In this paper, we investigate this issue by fine-tuning LLMs on three representative telecom datasets and show that safety degrades even for light telecom domain adaptation. To this end, we introduce TeleHarm, the first telecom-specific red-teaming benchmark, which we use alongside established Direct-Harm and HexPhi datasets to systematically assess harmful behavior. We further extend our analysis to publicly available TeleLLMs that were continually pre-trained on large telecom corpora, revealing that safety alignment is severely lacking, primarily due to the omission of safety-focused instruction tuning. To address these issues, we evaluate three realignment defenses: SafeInstruct, SafeLoRA, SafeMERGE. We show that, across all settings, the proposed defenses can effectively restore safety without compromising telecom task performance, leading to Safe teleCOMMunication (SafeCOMM) models. Our work serves as both a diagnostic study and practical guide for safety realignment in telecom-tuned LLMs, underscoring the need for safety-aware instruction and fine-tuning in the telecom domain.
@@ I Introduction and Motivation
Nonetheless, maintaining safety alignment is challenging when LLMs are further fine-tuned for downstream tasks. Recent work in [6] shows that a handful of malicious training examples can already jailbreak LLMs, causing them to comply with harmful user requests. In addition, the authors in [7] demonstrate that even conventionally benign fine-tuning can inadvertently compromise safety to a noticeable degree. Theoretical insights on this were later explored by investigating refusal directions and token-depth, suggesting that safety alignment is often shallow and easily disrupted [8, 9].
These findings confirm a problematic reality: LLMs may override their safety guardrails when adapted for new domains. Therefore, ensuring that telecom LLMs remain safe after fine-tuning is an important practical challenge. In this paper, we address this by investigating SFT of widely used Llama [10, 11] and Qwen [12] models on three publicly available telecom datasets: TeleQnA [13], TeleData [4], and TSpecLLM [14]. In addition, we examine publicly released TeleLLMs [4], which underwent continual pre-training using large telecom corpora. Our main goal is to confirm that even seemingly benign telecom datasets are not exempt from safety degradation, and to demonstrate that safety can be restored using lightweight methods. Our key contributions and findings are:
@@
    We introduce TeleHarm, the first telecom-specific red-teaming dataset, featuring harmful prompts from across the network stack and complementing general-purpose benchmarks such as DirectHarm [15] and HexPhi [7].
@@
    Supervised fine-tuning (SFT) on telecom data leads to noticeable safety degradation, as measured across DirectHarm, HexPhi, and TeleHarm benchmarks.
  3. 3.
    Continual pre-training (CPT) without safe instruction tuning leads to harmfulness ratios close to 90%. where TeleLLMs comply to most harmful user prompts.
@@ II-C Why Telecom Data May Be Prone to Safety Degradation
Telecom datasets often contain large volumes of tabular entries (e.g., 3GPP tables), long documents (e.g., standardization reports), and specialized protocols with complex formulas and algorithms. As shown in [21], such structured tabular and math-heavy data can amplify safety degradation by exposing LLMs to specialized jargon and formats that resemble harmful samples. Specifically, the authors demonstrate that benign data, particularly when formatted as lists, bullet points, or math problems, can significantly increase harmfulness after SFT. This occurs because these samples often share update directions in gradient space with known harmful data. Given the prevalence of these formats in telecom datasets, this raises an important open question: Does SFT or CPT with telecom data lead to similar or even worse safety degradation? Our goal is to address this question using open-source telecom datasets, investigating the importance of safety-aware data curation even in domains typically perceived as benign. In addition, our work provides guidance for practitioners unfamiliar with AI safety, showing how safe telecom LLMs can be trained with open-source frameworks without deep expertise in safety alignment.
@@ III-B SafeLoRA / SafeMERGE: Layer-Wise LoRA Adaptation
In most practical scenarios, SFT is implemented via parameter-efficient LoRA training [22], which introduces low-rank adapters into the model layers. For a weight matrix $W^{i}\in\mathbb{R}^{d\times k}$ in a transformer layer $i$, LoRA introduces two trainable matrices $A^{i}\in\mathbb{R}^{d\times r}$ and $B^{i}\in\mathbb{R}^{r\times k}$ (with $r\ll\min(d,k)$) such that the adapted weight becomes $W_{\text{LoRA}}^{i}=W^{i}+\Delta W^{i}=W^{i}+\gamma\cdot A^{i}B^{i}$, where $\gamma$ is a scaling factor. During fine-tuning, $W^{i}$ remains frozen while only $A^{i}$ and $B^{i}$ are trained. Both SafeLoRA and SafeMERGE selectively adapt only those LoRA layers that exhibit harmful behavior. To this end, they introduce a safety-aligned subspace $V^{i}$, computed as the difference between the weights of the base (unaligned) and instruct (aligned) version of the model. This subspace represents the safety alignment in the weight space and the projection $C^{i}$ onto it can be computed by:
@@
    $\Delta W_{\text{project}}^{i}=C^{i}\Delta W_{f}^{i}$ (5)
@@
    $\Delta W_{\mathrm{merge}}^{i}=\alpha\Delta W_{f}^{i}+(1-\alpha)\,\Delta W_{s}^{i}\ ,\ \alpha\in(0,1)$ (6)
@@ IV-A SFT Fine-Tuning on Telecom Datasets
We fine-tune three widely used instruction-tuned models: Llama-2-7B-Chat [10], Llama-3.1-8B-Instruct [11], and Qwen-2-7B-Instruct [12]. For datasets, we choose the QA-formatted benchmark datasets from TeleQnA [13] (8k samples), TeleData [4] (600k samples), and TSpecLLM [14] (80 samples), where we created 80/20 train-test splits, respectively. These datasets contain various telecom-specific questions drawn from standards, implementations, and engineering practice, often formatted as lists, tables, and complex mathematical formulas. The varying dataset sizes further allow us to analyze how the amount of telecom data impacts safety after SFT. All models are fine-tuned using Llama-Factory with an effective batch size of 32 and a learning rate of 1e-4 with linear scheduling. We train for 2 epochs on TeleData and for 5 epochs on TeleQnA and TSpecLLM. After SFT, we evaluate model performance on the test split by following the approach in [4] where we use Mixtral-8x7B-Instruct [23] as a judge to compare answers with ground truth responses. We compute the final accuracy as the ratio of correctly answered questions.
@@ IV-B Safety Evaluations
To assess safety, we generate responses on DirectHarm [15] and HexPhi [7], two general-purpose red-teaming datasets containing intentionally harmful prompts. In addition, we introduce TeleHarm [^1] Dataset available at: https://huggingface.co/aladinDJ/TeleHarm, a telecom-specific red-teaming benchmark of 125 harmful prompts carefully curated by domain experts across five categories: Physical/RF Layer, Access & Authentication, Core Network & Signaling, Infrastructure & OSS/BSS, and Privacy & Social Engineering. With TeleHarm, we target threat scenarios that are unique to communication systems and that span the entire network stack, providing, to the best of our knowledge, the first comprehensive telecom-specific benchmark that captures both high- and low-level risks, thereby complementing general-purpose red-teaming datasets. Following standard practice, we use a safety LLM judge (Llama-Guard-3-8B [11]) to evaluate model safety and report the overall harmfulness score as the proportion of generated responses flagged as unsafe.
@@ IV-D Open-Source CPT Telecom Models
  • SafeLoRA / SafeMERGE: we extract the LoRA layers from CPT models and apply the same safety projection and merging strategies. No additional training is required.
@@ V Results and Discussions
Qwen-2-7B-Instruct TeleQnA $(\uparrow)$ 45.80 65.60 64.80 64.10 65.20
TeleHarm $(\downarrow)$ 7.00 14.50 8.70 9.40 7.90
@@
TSpecLLM (80 samples) Llama-2-7B-Chat TSpecLLM $(\uparrow)$ 33.30 44.20 43.90 42.90 43.80
@@
Qwen-2-7B-Instruct TSpecLLM $(\uparrow)$ 12.50 28.30 28.00 27.70 28.10
@@
TABLE III: Task performance and harmfulness scores for Llama and Gemma TeleLLMs, evaluated on TeleData, TeleQnA, and TSpecLLM. The lack of safety samples during instruction tuning leads to extremely high harmfulness. SafeInstruct, SafeLoRA, and SafeMERGE defenses can effectively restore safety.
@@ V-A Telecom Task Utility after SFT and CPT
For CPT, task utility improves similarly between 10% and 15% on the TeleData benchmark for both Gemma and Llama-3 TeleLLMs, with similar trends observed for TeleQnA and TSpecLLM (see Table III). These results suggest that CPT, as performed by the authors in [4], effectively instills telecom knowledge, measured across diverse telecom benchmarks.
@@ V-B Harmfulness after SFT
For SFT, harmfulness increases noticeably, confirming that fine-tuning with telecom data degrades model safety. We observe the sharpest decline for Llama-2 on the larger TeleData dataset, which sees its DirectHarm (HexPhi) score rise from 5.00% (2.00%) to 36.70% (20.10%), followed by Qwen-2, whose original model is already the least safe. We see similar trends on our TeleHarm benchmark, with Llama-2 reaching 28%, Qwen-2 26%, and Llama-3.1 22%, confirming that harmfulness extends even to telecom-specific prompts. Scores on TeleQnA and TSpecLLM are comparable, suggesting a similar impact despite TSpecLLM being significantly smaller.
@@ V-D Safety Realignment
For SFT and CPT, SafeInstruct, SafeLoRA, and SafeMERGE significantly improve safety while maintaining telecom task utility. For Llama-2 on TeleData, either defense preserves accuracy around 38% while reducing harmfulness by up to 30% (15%) on DirectHarm (HexPhi) and by 25% on TeleHarm when using SafeMERGE, which generally provides the best safety-utility trade-off among the examined defenses, followed by SafeInstruct. For Llama-3 and Qwen-2, harmfulness can, in most cases, be reduced even below that of the original safety-tuned models. For CPT, all three defenses similarly reduce harmfulness from previously extreme levels to low double-digit scores with strong utility. We found thresholds $\tau$ around 0.6 to be optimal for both SafeLoRA and SafeMERGE (with $\alpha$ values of 0.7), such that only a small portion of LoRA layers need to be adapted. However, SafeMERGE requires tuning two parameters instead of one for SafeLoRA. SafeInstruct is the easiest to implement while requiring relatively few safety samples. We further added telecom-inspired safety refusals (e.g., for our base station example) but observed no notable gains during testing, suggesting that Bianchi’s data from [16] generalizes well, even for the telecom domain. Ultimately, these results show that either defense can reduce harmfulness without compromising utility, particularly on TeleHarm, which targets telecom threat scenarios across the network stack.
@@ V-E Per-Token KL Divergence between Safe and Unsafe Models
In Fig. 3, we illustrate the per-token KL divergence on unsafe TeleHarm prompts between fine-tuned models and their unaligned base counterparts, capturing how quickly the distributions drift toward unsafe behavior. This provides a token-level-proof that SFT with telecom data consistently degrades safety. While all three telecom models exhibit high KL divergence in the first few tokens, their distributions quickly converge toward the unsafe base model, indicating that alignment is only a few tokens deep. As a result, models may eventually respond to harmful prompts despite an initial refusal (e.g., “I cannot, but …”). Degradation differs across datasets: TSpecLLM diverges most, followed by TeleQnA and TeleData, likely because its smaller training set keeps it initially closer to the safe instruct model. This shows that safety erosion depends on the number of potentially harmful telecom samples during training. We also include the divergence between the SafeMERGE variant and its safety-aligned Llama-3 instruct model, showing that SafeMERGE closely matches the instruct model, thereby restoring safety alignment from the first tokens.
@@ VI Conclusion
In this paper, we studied the impact of fine-tuning LLMs to the telecom domain on model safety, demonstrating that both SFT and CPT can significantly degrade safety alignment. To benchmark this problem, we introduced TeleHarm, a novel telecom-specific red-teaming dataset, and investigated model safety after fine-tuning on three representative SFT datasets as well as two publicly available TeleLLMs continually pre-trained on large-scale telecom corpora. Our findings show that incorporating safety-aligned instruction tuning is necessary, as technical telecom data can inadvertently resurface or amplify harmful behaviors present in the base model. We further demonstrated that lightweight, open-source safety realignment methods can effectively restore safety while preserving strong telecom task utility. Our study thus underscores a key takeaway: safety alignment should not be an afterthought in the development of telecom LLMs and can be addressed either early or post-hoc with little effort and substantial impact.

### v3

@@ Abstract
Fine-tuning large language models (LLMs) on telecom datasets is a common practice to adapt general-purpose models to the telecom domain. However, little attention has been paid to how this process may compromise model safety. Recent research has shown that even benign fine-tuning can degrade the safety alignment of LLMs, causing them to respond to harmful or unethical user queries. In this paper, we investigate this issue by fine-tuning LLMs on three representative telecom datasets and show that safety degrades even for light telecom domain adaptation. To this end, we introduce TeleHarm, the first telecom-specific red-teaming benchmark, which we use alongside established DirectHarm and HexPhi datasets to systematically assess harmful behavior. We further extend our analysis to publicly available TeleLLMs that were continually pre-trained on large telecom corpora, revealing that safety alignment is severely lacking, primarily due to the omission of safety-focused instruction tuning. To address these issues, we evaluate three realignment defenses: SafeInstruct, SafeLoRA, SafeMERGE. We show that, across all settings, the proposed defenses can effectively restore safety without compromising telecom task performance, leading to Safe teleCOMMunication (SafeCOMM) models. Our work serves as both a diagnostic study and practical guide for safety realignment in telecom-tuned LLMs, underscoring the need for safety-aware instruction and fine-tuning in the telecom domain.
@@ I Introduction and Motivation
Nonetheless, maintaining safety alignment is challenging when LLMs are further fine-tuned for downstream tasks. Recent work in [6] shows that a handful of malicious training examples can already jailbreak LLMs, causing them to comply with harmful user requests. In addition, the authors in [7] demonstrate that even conventionally benign fine-tuning can inadvertently compromise safety to a noticeable degree. Subsequent work provides mechanistic explanations by investigating refusal directions and token-depth, suggesting that safety alignment is often shallow and can be easily disrupted [8, 9].
These findings confirm a problematic reality: LLMs may override their safety guardrails when adapted for new out-of-distribution domains. Therefore, ensuring that telecom LLMs remain safe after fine-tuning is an important practical challenge. In this paper, we address this by investigating SFT of widely used Llama [10, 11] and Qwen [12] models on three publicly available telecom datasets: TeleQnA [13], TeleData [4], and TSpecLLM [14]. In addition, we examine publicly released TeleLLMs [4], which underwent continual pre-training using large telecom corpora. Our main goal is to confirm that even seemingly benign telecom datasets are not exempt from safety degradation and to demonstrate that safety can be restored using lightweight methods. To this end, we introduce TeleHarm, the first telecom-specific red-teaming dataset, featuring harmful prompts from across the network stack, thus complementing general-purpose benchmarks such as DirectHarm [15] and HexPhi [7]. Our key findings are:
@@
    Supervised fine-tuning (SFT) on telecom data leads to noticeable safety degradation as measured across Direct-Harm, HexPhi, and TeleHarm benchmarks.
@@
    Continual pre-training (CPT) without safe instruction tuning leads to harmfulness ratios close to 90%, such that TeleLLMs comply with most harmful user prompts.
@@ II-C Why Telecom Data May Be Prone to Safety Degradation
Telecom datasets often contain large volumes of tabular entries (e.g., 3GPP tables), long documents (e.g., standardization reports), and specialized protocols with complex formulas and algorithms. As shown in [21], such structured tabular and math-heavy data, particularly when formatted as lists, bullet points, or math equations, can amplify safety degradation by exposing LLMs to specialized jargon and formats that resemble harmful samples. This occurs because these samples often share update directions in gradient space with known harmful data. Given the prevalence of these formats in telecom datasets, this raises an important open question: Does SFT or CPT with telecom data lead to similar or even worse safety degradation? Our goal is to address this question using open-source telecom datasets, investigating the importance of safety-aware data curation even in domains typically perceived as benign. While other factors such as dataset scale and training dynamics may further influence safety, our results show that format-level properties alone are sufficient to trigger degradation. Importantly, our work provides actionable guidance for practitioners unfamiliar with AI safety by demonstrating how safe telecom LLMs can be trained with open-source frameworks without deep expertise in safety alignment techniques.
@@ III-B SafeLoRA / SafeMERGE: Layer-Wise LoRA Adaptation
In most practical scenarios, SFT is implemented via parameter-efficient LoRA training [22], which introduces low-rank adapters into the model layers. For a weight matrix $W^{i}\in\mathbb{R}^{d\times k}$ in a transformer layer $i$, LoRA introduces two trainable matrices $A^{i}\in\mathbb{R}^{d\times r}$ and $B^{i}\in\mathbb{R}^{r\times k}$ (with $r\ll\min(d,k)$) such that the adapted weight becomes $W_{\text{LoRA}}^{i}=W^{i}+\Delta W^{i}=W^{i}+\gamma\cdot A^{i}B^{i}$, where $\gamma$ is a scaling factor. During fine-tuning, $W^{i}$ remains frozen while only $A^{i}$ and $B^{i}$ are trained. Both SafeLoRA and SafeMERGE selectively adapt only those LoRA layers that exhibit harmful behavior. To this end, they introduce a safety-aligned subspace $V^{i}$, computed as the difference between the weights of the base (unaligned) and instruct (aligned) version of the model. This subspace represents the safety alignment in the weight space and the projection $C^{i}$ onto it can be computed as:
@@
    $\Delta W_{\text{project}}^{i}=C^{i}\Delta W_{f}^{i}\ .$ (5)
@@
    $\Delta W_{\mathrm{merge}}^{i}=\alpha\Delta W_{f}^{i}+(1-\alpha)\,\Delta W_{s}^{i}\ ,\ \alpha\in(0,1)\ .$ (6)
@@ IV-A SFT Fine-Tuning on Telecom Datasets
We fine-tune three widely used instruction-tuned models: Llama-2-7B-Chat [10], Llama-3.1-8B-Instruct [11], and Qwen-2-7B-Instruct [12]. For datasets, we choose the QA-formatted benchmark datasets from TeleQnA [13] (8k samples), TeleData [4] (600k samples), and TSpecLLM [14] (80 samples), where we created 80/20 train-test splits, respectively. These datasets contain various telecom-specific questions drawn from standards, implementations, and engineering practice, often formatted as lists, tables, and complex mathematical formulas. The varying dataset sizes further allow us to analyze how the amount of telecom data impacts safety after SFT. All models are fine-tuned using Llama-Factory with an effective batch size of 32 and a learning rate of 1e-4 with linear scheduling. We train for 2 epochs on TeleData and for 5 epochs on TeleQnA and TSpecLLM. After SFT, we evaluate model performance on the test split by following the approach in [4], where we use Mixtral-8x7B-Instruct [23] as a judge to compare answers with ground truth responses. We compute the final accuracy as the ratio of correctly answered questions.
@@ IV-B Safety Evaluations
To assess safety, we generate responses on DirectHarm [15] and HexPhi [7], two general-purpose red-teaming datasets containing intentionally harmful prompts. In addition, we introduce TeleHarm [^1] Dataset available at: https://huggingface.co/datasets/aladinDJ/TeleHarm, a telecom-specific red-teaming benchmark of 125 harmful prompts carefully curated by domain experts across five categories: 1) Physical & RF Layer, 2) Access & Authentication, 3) Core Network & Signaling, 4) Infrastructure & OSS/BSS, and 5) Privacy & Social Engineering. With TeleHarm, we target threat scenarios that are unique to communication systems and that span the entire network stack, thus providing, to the best of our knowledge, the first comprehensive telecom-specific benchmark that captures both high- and low-level risks, thereby complementing general-purpose red-teaming datasets. Following standard practice, we use a safety LLM judge (Llama-Guard-3-8B [11]) to evaluate model safety and report the overall harmfulness score as the proportion of generated responses flagged as unsafe.
@@ IV-D Open-Source CPT Telecom Models
  • SafeLoRA / SafeMERGE: we extract LoRA layers from the CPT models and apply the same safety projection and merging strategies. No additional training is required.
@@ V Results and Discussions
Qwen-2-7B-Instruct TeleQnA $(\uparrow)$ 45.80 65.60 64.80 64.10 65.60
TeleHarm $(\downarrow)$ 7.00 14.50 8.70 9.40 6.90
@@
TSpecLLM (80 samples) Llama-2-7B-Chat TSpecLLM $(\uparrow)$ 33.30 44.20 43.90 42.90 43.90
@@
Qwen-2-7B-Instruct TSpecLLM $(\uparrow)$ 12.50 28.30 28.00 27.70 28.30
@@
TABLE III: Task performance and harmfulness scores for Llama and Gemma TeleLLMs, evaluated on TeleData, TeleQnA, and TSpecLLM. The lack of safety samples during instruction tuning leads to extremely high harmfulness. SafeInstruct, SafeLoRA, and SafeMERGE defenses can significantly improve safety.
@@ V-A Telecom Task Utility after SFT and CPT
For CPT, task utility improves similarly between 10% and 15% on the TeleData benchmark for both Gemma and Llama TeleLLMs, with similar trends observed for TeleQnA and TSpecLLM (see Table III). These results suggest that CPT, as performed by the authors in [4], effectively instills telecom knowledge, measured across diverse telecom benchmarks.
@@ V-B Harmfulness after SFT
For SFT, harmfulness increases noticeably, confirming that fine-tuning with telecom data degrades model safety. We observe the sharpest decline for Llama-2 on the larger TeleData dataset, which sees its DirectHarm (HexPhi) score rise from 5.00% (2.00%) to 36.70% (20.10%), followed by Qwen-2, whose original model is already the least safe. We see similar trends on our TeleHarm benchmark, with Llama-2 reaching 28%, Qwen-2 26%, and Llama-3.1 22%, confirming that harmfulness extends to telecom-specific prompts. Scores on TeleQnA and TSpecLLM are comparable, suggesting a similar impact despite TSpecLLM being significantly smaller.
@@ V-D Safety Realignment
For SFT, SafeInstruct, SafeLoRA, and SafeMERGE significantly improve safety while maintaining telecom task utility. For Llama-2 on TeleData, either defense preserves accuracy around 38% while reducing harmfulness by up to 30% (15%) on DirectHarm (HexPhi) and by 25% on TeleHarm when using SafeMERGE, which generally provides the best safety-utility trade-off. For Llama-3 and Qwen-2, harmfulness can, in most cases, be reduced even below that of the original models. For CPT, all three defenses similarly reduce harmfulness from previously extreme levels to low double-digit scores with strong utility. We found thresholds $\tau$ around 0.6 to be optimal for both SafeLoRA and SafeMERGE (with $\alpha$ values of 0.7), such that only a small portion of LoRA layers need to be adapted. However, SafeMERGE requires tuning two parameters instead of one for SafeLoRA. SafeInstruct is the easiest to implement while requiring relatively few safety samples. We further added telecom-inspired safety refusals (e.g., for our base station example) but observed no notable gains during testing, suggesting that Bianchi’s data [16] generalizes well to the telecom domain. These results show that either defense can reduce harmfulness without compromising utility, particularly on TeleHarm, which targets threat scenarios across the stack.
@@ V-E Per-Token KL Divergence between Safe and Unsafe Models
In Fig. 3, we illustrate the per-token KL divergence on unsafe TeleHarm prompts between fine-tuned models and their unaligned base counterparts. While all three telecom models exhibit high KL divergence in the first few tokens, their distributions quickly converge toward the unsafe base model, indicating that alignment is only a few tokens deep. As a result, models may eventually respond to harmful prompts despite an initial refusal (e.g., “I cannot, but …”). This provides a token-level-proof that SFT with telecom data consistently degrades safety. Degradation further differs across datasets: TSpecLLM diverges most, followed by TeleQnA and TeleData, likely because its smaller training set keeps it initially closer to the safe instruct model. This shows that safety erosion may depend on the number of potentially harmful telecom samples during training. We also include the divergence between the SafeMERGE variant and its safety-aligned Llama-3 instruct model, showing that SafeMERGE closely matches the instruct model, thereby restoring safety alignment from the first tokens.
@@ VI Conclusion
In this paper, we studied how fine-tuning LLMs to the telecom domain impacts model safety. To benchmark this problem, we introduced TeleHarm, a telecom-specific red-teaming dataset, and investigated model safety after fine-tuning on three representative SFT datasets as well as on two publicly available TeleLLMs continually pre-trained on large telecom corpora. Our findings show that incorporating safety-aligned instruction tuning is necessary, as technical telecom data can inadvertently resurface or amplify harmful behaviors present in the base model. We further demonstrated that lightweight safety realignment methods can effectively restore safety while preserving strong task utility. Our study thus underscores a key takeaway: safety alignment should not be an afterthought in the development of telecom LLMs and can be addressed either early or post-hoc with little effort and substantial impact.